A truck travels for 39.1 minutes at a speed of 55.5 km/h and then for 49.7 minutes at a speed of 98.9 km/h. What is the average speed of the truck?
Answer #1

given that

time t1 - 39.1 min = 39.1/60 = 0.652 hr

v1 = 55.5 km/h

t2 = 49.7 min = 0.828 hr

v2 = 98.9 km/h

distance during the time t1 is

d1 = v1*t1 = 55.5*0.652 = 36.186 km

during time t2

d2 = v2*t2 = 98.9*0.828 = 81.89 km

average speed v avg = d1+d2/t1+t2

v avg = (36.186+81.89)/(0.652+0.828) = 79.78 km/h